Calculate the price elasticity of demand using the mid-point Method from the following information.

Price

20

30

40

Quantity

90

60

45

Would your result be different if instead you consider the total revenue approach to calculate the price elasticity of demand?

b. Show that a minimum wage set above the market clearing wage rate causes excess supply of labour. Why would a government impose minimum wage despite knowing its effect on unemployment?

c. What is effect on producer surplus if taste changes in favor of organic food. Explain it diagrammatically. Will it affect producer surplus in the market for organic food?


1
Expert's answer
2021-03-31T11:57:00-0400

a. The price elasticity of demand using the mid-point method is:

"Ed = \frac{60 - 90} {30 - 20} ×\frac{30 + 20} {60 + 90} = -1."

So, the demand is unit-elastic at this point.

The result would be the same if instead you consider the total revenue approach to calculate the price elasticity of demand.

b. If a minimum wage set above the market clearing wage rate, then the quantity of labour demanded will be lower than the quantity of labour suppied, as a result excess supply of labour will occur.

Governments use minimum wage laws to ensure a basic quality of life among all citizens within its borders. These laws attempt to improve an individual's position in the economic income brackets. Rather than have copious amounts of underpaid or poor citizens, minimum wage laws seek a level of economic equality.

c. If taste changes in favor of organic food, then the demand for organic food will increase, and the producer surplus will increase too.